OpenStudy (anonymous):

Solve this problem. Brianna paid $18.89 for a pair of sandals. She bought a pair of hiking boots that cost $12 more than her sandals. How much did Brianna pay for the two pairs of shoes? How can estimation be used to make sure the answer is reasonable? A. Answer: Brianna paid $42.89. Estimation: $19 + $12 = $31 $31 + $12 = $43 Brianna paid about $43 for the shoes. B. Answer: Brianna paid $25.78. Estimation: $19 – $12 = $7 $19 + $7 = $26 Brianna paid about $26 for the shoes.

OpenStudy (anonymous):

C. Answer: Brianna paid $49.78. Estimation: $19 + $12 = $31 $31 + $19 = $50 Brianna paid about $50 for the shoes.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@Mahoganie.Carson

OpenStudy (mahoganie.carson):

I think it's A

OpenStudy (helpkay):

yes i would say it is c the fist pair of show are 19$ the other pair of shoes are 31$ because it said the price of the first pair of shoes plush 12$ 19+13=31 then it say what it the cost of both shoes 31+19=50 so I would have to say C

OpenStudy (anonymous):

2 shoes cost 19 for sandals 31 for boots = 50 or to be exact 49.78 So I say C

OpenStudy (helpkay):

19+12=31 not 19+13

OpenStudy (mahoganie.carson):

Yeah @countsrnc I think she's right @helpkay

OpenStudy (mahoganie.carson):

What i was telling her that your right @helpkay

OpenStudy (helpkay):

please fan and medal :)

OpenStudy (mahoganie.carson):

@countsrnc tell us if you got it right later :)